Find the minimum kinetic energy (in MeV) in the lab system needed by an alpha particle to cause


the reaction:


7𝑁


14 + 𝛼 β†’ 8𝑂


17 + 1𝐻


1


Given that: m(


14N) = 14.00307 u; m(Ξ±) 4.003879 u; m(


1H) = 1.00783 u; m(


17O) = 17.0013 u;


1 𝑒 = 931.5 𝑀𝑒𝑉/𝑐


2

1
Expert's answer
2022-01-10T09:09:36-0500

"E=\Delta mc^2=(14.00307+4.003879-1.00783-17.0013)\cdot 931 .5=2.03~MeV."